OpenStudy (anonymous):

How maximum f (x)=x^3 - 12x + 2

OpenStudy (aum):

Find the derivative, equate it to zero and solve for x. Find the second derivative and test each critical point found in step one. For maximum, f''(c) < 0.

OpenStudy (anonymous):

I don't understand Please Explain to me

OpenStudy (aum):

Do you know how to take the derivative? If yes, can you find the derivative of f(x)?

OpenStudy (anonymous):

No l can't

OpenStudy (anonymous):

Please halp me

OpenStudy (aum):

\[ f (x)=x^3 - 12x + 2 \\ f'(x) = 3x^2 - 12 = 0 \\ 3x^2 = 12 \\ x^2 = 4 \\ x = -2 \text{ or } x = 2 \\ f''(x) = 6x \\ f''(-2) = -12 \\ \text{Since f''(-2) < 0 x = -2 gives a maximum.}\\ \text{The maximum value is } f(-2) = (-2)^3 - 12(-2) + 2 = -8 + 24 + 2 = 18 \]

OpenStudy (anonymous):

Thank you Can you answer to another question?

OpenStudy (anonymous):

X^3-x^2+1

OpenStudy (aum):

Follow the method shown above and apply the same method for the second problem.

OpenStudy (anonymous):

I could not be resolved Please Explain it